OpenStudy (anonymous):

Help!!!!!!!! Solve for x : log(base3)(4.3^x-1) = 2x+1

OpenStudy (anonymous):

help quickly friends!!!!!!!

ganeshie8 (ganeshie8):

\[\log_{3}(4 * 3^{(x-1)}) = 2x+1 \]

ganeshie8 (ganeshie8):

you equations looks like above ?

OpenStudy (anonymous):

no it is a decimal point.

OpenStudy (anonymous):

4.3(power-x)-1

OpenStudy (anonymous):

is it 4 into 3 to power (x-1) or 4*3 to power (x) -1 i.e. 4*3^(x-1) or 4*3^(x) -1

ganeshie8 (ganeshie8):

\[\log_{3}(4.3^{x}-1) = 2x-1 \]

OpenStudy (anonymous):

it is 4.3 to the power x

ganeshie8 (ganeshie8):

ok it looks like above is it ?

OpenStudy (anonymous):

yes ganesh's equation is right

ganeshie8 (ganeshie8):

this doesnt look simple to me to solve... leaving it to himanshu.. :)

OpenStudy (anonymous):

solve 4.3^x-1=3^(2x+1) to get x

OpenStudy (anonymous):

log\[\log_{3}(4.3^{x}-1)=2x+1\]

OpenStudy (shubhamsrg):

it'd require the knowledge of higher mathematics (interpolation in this case) i'd guess//

OpenStudy (anonymous):

first u have to cancel the log.. to do this write (2x+1) as log(base 3) 3^(2x+1)

OpenStudy (shubhamsrg):

unless its 4 * (3^x) which makes it quad eqn..

OpenStudy (anonymous):

i want an answer

OpenStudy (shubhamsrg):

it must be 4 * (3^x) buddy,,am damn sure..(though i may be wrong :P)

OpenStudy (anonymous):

ok solve it as 4*(3^x)....................

OpenStudy (shubhamsrg):

let 3^x = y so you'll have 3y^2 - 4y + 1 =0 = (3y-1)(y-1) => y =1 and 1/3 so x =0 and -1

OpenStudy (anonymous):

x=0 and x=-1

OpenStudy (anonymous):

thanks himanshu and shubham

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

one more thiing when u solve for x for these question which contain log try to verify for all values of x u have found..there may be a case u get negative in a log for some x. u have to discasr that value of x

OpenStudy (anonymous):

*discard

OpenStudy (anonymous):

ok thanks

OpenStudy (shubhamsrg):

hmmn,, :)